For Stats- If A:(1,3,5,7) and B: (1,2,3,4,5), find P(B/A)?

Answer by stanbon(75887)

For Stats- If A:(1,3,5,7) and B: (1,2,3,4,5), find P(B/A)?
----
P(B|A) = P(A and B)/P(B) = P{1,3,5}/P{1,2,3,4,5) = 3/5
====================
Cheers,
Stan H.
================